OpenStudy (anonymous):

Giving Medals ((&)) A Fan Below are the steps to solve an equation: Step 1: |x - 2| + 3 = 7 Step 2: |x - 2| = 7 - 3 Step 3: |x - 2| = 4 Which of the following is a correct next step to solve the equation? A) x + 2 = -4 B) -x - 2 = 4 C) x + 2 = 4 D) x - 2 = -4

OpenStudy (anonymous):

have you learned about absolute values?

OpenStudy (anonymous):

yeah but algebra is not my strongest area

OpenStudy (the_fizicx99):

D

OpenStudy (anonymous):

Thats okay :)) To clear the absolute-value bars, I must split the equation into its two possible two cases, one case for each sign.

OpenStudy (anonymous):

So in this case your answer would be C, I think

OpenStudy (help!!!!):

x-2=4 or x-2=-4 solve

OpenStudy (anonymous):

Thx @tHe_FiZiCx99 how @Melodysim

OpenStudy (the_fizicx99):

Yw

OpenStudy (help!!!!):

No it would be D

OpenStudy (anonymous):

I said you have to split the equation into its two possible two cases, one case for each sign.

OpenStudy (anonymous):

so one with positive and other with negative

OpenStudy (anonymous):

http://www.purplemath.com/modules/solveabs.htm Go check this out :) Read it and try other questions

OpenStudy (help!!!!):

The the two outcomes should be x-2=4 OR x-2=-4

OpenStudy (anonymous):

So that would make it D ?

OpenStudy (help!!!!):

YES the answer is D

OpenStudy (anonymous):

Wait lol I said C? sorry for the type o wow lmao

OpenStudy (anonymous):

It is D sorry about that :(

OpenStudy (anonymous):

Okay thanks but can you help me with one more question @HELP!!!! @Melodysim btw its okay :)

OpenStudy (help!!!!):

yeah but a medal would be appreciated

OpenStudy (anonymous):

Solve 4(x + 1) = 4x + 10. A) 0 B)6 C)All real numbers D)No solution and ok how do you give a medal ?

OpenStudy (help!!!!):

just click best response on me

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Now can you help me with the question ? @HELP!!!!

OpenStudy (help!!!!):

it would be no solution because \[4\neq-6\] and you know how to solve this right?

OpenStudy (anonymous):

Yes thanks alot !